OpenStudy (anonymous):

find the dimensions of a rectangle whose length is 5 less than twice its width and whose area is 63 sq. units.

OpenStudy (amistre64):

we know that: Area = 63 = lw; where w=w and l=2w-5 63 = (2w-5)w 63 = 2w^2-5w 63 = 2(w^2-5w/2) 63 + 50/16 = 2(w^2-5w/2 + 25/16) 63 + 50/16 = 2(w-5/4)^2 (63 + 50/16)/2 = (w-5/4)^2 sqrt((63 + 50/16)/2) = w-5/4 sqrt((63 + 50/16)/2) + 5/4 = w = 7, so maybe :)

OpenStudy (amistre64):

9*7 = 63 so heres hopin ... l =2(7) - 5 = 14 - 5 =9 ..........yay!!